# Optimizing hypergraph product codes with random walks, simulated annealing and reinforcement learning

### v1

@@ II-B3 Cost function
implying the existence of a logical error supported within the erasure. In that case, there exist two errors supported on the erasure, with the same syndrome, but differing by a nontrivial logical operator. In the following, we define as logical error rate the probability that a random erasure (with erasure parameter $p$) admits such a $v$, and therefore cannot be corrected by a maximum-likelihood decoder.
@@ II-C1 Plain exploration
The simplest optimization strategy is a naive exploration of the state space, navigated through by the application of actions. The plain exploration starts from a given initial state, and performs a random walk of length $L$. At each step, the algorithm computes the cost function of the current state, as well that the cost function of $N-1$ randomly chosen neighbors of the current state. In total, the plain exploration strategy computes $NL$ times the cost function.
@@ III Results and discussion
Fig. 1 shows the performance over the erasure channel with maximum likelihood decoding of the codes obtained from the different optimization strategies [^1] Note that our criterion based on the existence of logical errors supported within the erasure is stricter than the one presented in Connolly et al., explaining the discrepancy with the results of Ref. [24].. While each code family was optimized for a single value of the erasure channel (indicated in Table I), the simulation results show that the optimized codes with length 625 and 1600 outperform the initial ones over a whole range of parameters, and in particular in the regime of low logical error rate.
@@
Figure 1: Comparison of the performance under the erasure channel of the original code from [24] (PEG), with the best codes obtained through plain exploration (PE), simulated annealing (SA) and projective simulation (PS), estimated with $5\cdot 10^{5}$, $10^{5}$ and $5\cdot 10^{4}$ Monte Carlo samples for the respective code lengths.

### v3

@@ II-B3 Cost function
implying the existence of a logical error supported within the erasure. In that case, the erasure hides syndrome-compatible corrections that differ by a nontrivial logical operator. For such verification one must solve some linear systems, hence the cost function has $O(N^{3})$ complexity. In the following, we define as logical error rate the probability that a random erasure (with erasure parameter $p$) admits such a $v$, and therefore cannot be corrected by a maximum-likelihood decoder.
@@ II-C1 Plain exploration
The simplest optimization strategy is a naive exploration of the state space, navigated through by the application of actions. The plain exploration starts from a given initial state, and performs a random walk of length $L$. At each step, the algorithm computes the cost function of the current state, as well as the cost function of $N-1$ randomly chosen neighbors, that is, states resulting from the application of an action to the current state. In total, the plain exploration strategy computes the cost function $NL$ times.
@@ III Results and discussion
Fig. 1 shows the performance over the erasure channel with maximum likelihood decoding of the codes obtained from the different optimization strategies. While each code family was optimized for a single value of the erasure channel (indicated in Table I), the simulation results show that the optimized codes with length 625 and 1600 outperform the initial ones over a whole range of parameters, and in particular in the regime of low logical error rate. This is explained by their resulting code distances, annotated in the plots.
@@
Figure 1: Comparison of the performance under the erasure channel of the original code from [24] (PEG), with the best codes obtained through plain exploration (PE), simulated annealing (SA) and projective simulation (PS), estimated with $5\cdot 10^{5}$, $10^{5}$ and $5\cdot 10^{4}$ Monte Carlo samples for the respective code lengths. All PEG codes have girth $6$, whereas all optimized codes have girth $4$, which suggests that maximizing girth is not a good heuristics for improving performance.